In a group of 125 computer users, 50 of them have a GPU installed in their systems, 30 of them have SSD storage, and 15 have both. If a computer user chosen at random has a GPU, what is the probability he/she also has an SSD?


1
Expert's answer
2021-04-19T06:42:09-0400

Given,

Total computers "n(u)=125"

Number of computer having GPU installed in their systems "n(A)=50\Rightarrow P(A)=\dfrac{50}{125}=0.4"


Number of computer having ssd storage "n(B)=30\Rightarrow P(B)=\dfrac{30}{125}=0.24"


Number of computer having ssd storage as well as GPU


"n(A\cap B)=15\Rightarrow P(A\cap B)=\dfrac{15}{125}=0.12"



Required Probability "=\dfrac{P(A\cap B)}{P(A)}=\dfrac{0.12}{0.4}=0.3"